• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Michaels/Gayou

    Log entry time 11:18:19 on January13,2005

    Entry number 139172

    This entry is a followup to: 139022

    Followups:

    keyword=Trigger restore, tests

    We restored the trigger for the production running with protons on
    R-HRS, i.e. the "R-slow" setup. We did the change procedure in halog
    139022, modified slightly so that the true coinc time (TC) appears
    in the middle of the dynamic range of the TDC in the pulser setup.
    The slight modification affects the values of Ldelay1 and Rdelay2.

    Next, we test this with pulser setup by putting N nsec on L-HRS and
    subsequently N nsec on R-HRS in separate runs, to show that TC moves
    the expected way and that there is no delta-function goofiness like
    described in halog 139024 which affected the "Lslow" setup.

    CODA Run .... Delays ....... Coince Time (TC) measured on L,R
    ............. on pulser ..... ctimel ..... ctimer (raw channels)

    1777 ......... 0 relative. ... 1919 ..... 2068 (recall 1=0.05nsec)
    1778 ......... +30 to LHRS ... 1322 ..... 2684 (change = 600 = 30nsec)
    1779 ......... +30 to RHRS ... 2497 ..... 1468
    1780 ......... +60 to RHRS ... 2894(!) .. 850 ... problem w/cimel
    1781 ......... +60 to LHRS ... 707 ...... 3322

    All these are ok except ctimel for run 1780. But this is understood
    from my notes because the width of T3 was 74 nsec, and as T1 slides
    late it can make T5 over this interval but the retiming on LHRS has
    only a 47 nsec gap between L1A and RT. So, we reduce T3 width to
    40 nsec (must be less than 47). Do another run with RHRS by 45 nsec
    which is at the edge of the overlap (verified on scope):

    1782 ....... +45 to RHRS .... 2795 ..... 1161 ... Good !

    Final run to check the nominal setup

    1783 ...... 0 relative .... 1918 ..... 2072

    Finally, we remove the pulser. Test completed.